### 06 Mar 2024 v2

# Less is More: Hop-Wise Graph Attention for Scalable and Generalizable Learning on Circuits

Abstract
While graph neural networks (GNNs) have gained popularity for learning circuit representations in various electronic design automation (EDA) tasks, they face challenges in scalability when applied to large graphs and exhibit limited generalizability to new designs. These limitations make them less practical for addressing large-scale, complex circuit problems. In this work we propose HOGA, a novel attention-based model for learning circuit representations in a scalable and generalizable manner. HOGA first computes hop-wise features per node prior to model training. Subsequently, the hop-wise features are solely used to produce node representations through a gated self-attention module, which adaptively learns important features among different hops without involving the graph topology. As a result, HOGA is adaptive to various structures across different circuits and can be efficiently trained in a distributed manner. To demonstrate the efficacy of HOGA, we consider two representative EDA tasks: quality of results (QoR) prediction and functional reasoning. Our experimental results indicate that (1) HOGA reduces estimation error over conventional GNNs by $46.76\%$ for predicting QoR after logic synthesis; (2) HOGA improves $10.0\%$ reasoning accuracy over GNNs for identifying functional blocks on unseen gate-level netlists after complex technology mapping; (3) The training time for HOGA almost linearly decreases with an increase in computing resources. Source code of HOGA is freely available at: github.com/cornell-zhang/HOGA.
I Introduction
Recent years have seen a surge of interest in machine learning (ML) for electronic design automation (EDA), which holds great potential in achieving faster design closure and minimizing the need for extensive human supervision [10]. In particular, graph neural networks (GNNs) have become increasingly popular in the EDA community due to their ability to encode graph-structured data such as gate-level netlists into compact representations, which can be used for a multitude of downstream EDA applications, including quality of results (QoR) prediction and functional reasoning [13, 18].
However, scaling GNN training to large graphs is a notoriously challenging problem, which poses a serious concern on the practical benefit of GNNs on large-scale EDA problems. On the one hand, unlike common datasets on social networks and molecular graphs, which consist of either a few large graphs or a large number of small graphs, the circuit datasets may contain numerous graphs with each graph containing hundreds of thousands of nodes. For instance, the OpenABC-D benchmark provides $870$k gate-level netlists, where each netlist consists of up to $240$k logic gates [6]. Thus, training GNNs on such a large-scale circuit dataset is even more challenging than other graph-based applications. On the other hand, the mainstream GNN models are built upon the conventional message-passing paradigm among nodes, which learns node representations through a recursive node-wise aggregation scheme shown in Figure 1(b). As a consequence, it is nontrivial to perform efficient distributed GNN training due to the node dependencies in a graph structure.
Apart from the scalability challenge, it is also underexplored how to make GNNs generalizable across different circuit designs. Although there are many customized GNNs previously proposed for various EDA applications, their model backbones mainly follow classic GNNs such as GCN [11] and GraphSAGE [9], which are not necessarily suitable for circuit problems. Consider a task of identifying functional blocks within circuits [18]. As distinct functional blocks may have different depths, the number of hops to be considered varies across nodes, which cannot be easily captured by common GNNs. Moreover, the high-order structures of functional blocks are also important yet ignored by the aforementioned GNN models. As a result, existing GNNs for EDA tasks often struggle to learn the intrinsic and critical information from complex circuit graphs, resulting in limited generalizability to unseen designs.
Figure 1: Comparison of HOGA and conventional GNNs — (a) An example graph for illustration; (b) Computation graph of GNNs; (c) Computation graph of our proposed approach, HOGA.
In literature, improving the scalability and enhancing the generalizability of GNNs on circuits are largely viewed as two orthogonal directions and seldomly explored together. While there is a rich body of prior studies focusing on scaling GNNs to large graphs with distributed training, they either incur massive communication overhead among GPUs, or rely on heuristic graph sampling algorithms that may lose critical structural information [14], which in turn degrades the model generalizability. In the context of learning generalizable GNNs for circuit problems, Wang et al. [17] leveraged graph contrastive learning to pretrain GNNs and adopted multiple neighbor aggregation functions to learn the functionalities of logic gates, which comes with additional computational costs and thus worsens the model scalability.
Motivated by the limitations of conventional GNNs on circuit designs, we propose a novel hop-wise attention approach, named HOGA, to improve both the scalability and generalizability of circuit representation learning. As shown in Figure 1(c), HOGA adopts a hop-wise aggregation scheme, which precomputes the hop-wise features and only uses those features to learn node representations through a gated self-attention module. Since the final representation per node depends solely on its own hop-wise features, there are no dependencies between different nodes during training, making it easy to scale HOGA training in a distributed manner. Moreover, the gated self-attention module enables HOGA to adaptively aggregate important and high-order features from different hops per node, rendering it generalizable across different circuit designs.
It is noteworthy that HOGA is a flexible approach for learning circuit representations, which can be integrated with other customization techniques previously proposed for various downstream EDA tasks. To demonstrate the viability and flexibility of HOGA, we consider two representative circuit problems: (1) QoR prediction – We focus on predicting the optimized gate count after logic synthesis on the OpenABC-D benchmark, which is generated from various circuit designs as well as synthesis recipes, and is one of the largest open-sourced circuit datasets; (2) Functional reasoning – We follow the most challenging setting in Gamora [18] by identifying functional blocks on gate-level netlists after technology mapping. Our experiments show that HOGA not only outperforms prior GNNs on unseen designs in both applications, but also scales comfortably to large-scale circuits. We summarize our main technical contributions as follows:
$\bullet$ To the best of our knowledge, we are the first to introduce a scalable and generalizable model for circuit representation learning, which is achieved by a coarse-grained message-passing scheme based on hop-wise attention.
$\bullet$ By precomputing hop-wise features, HOGA avoids recursive neighbor aggregation during training. This enables HOGA to learn each node representation independently and facilitates massive parallelization for distributed training. As a result, the training time of HOGA almost linearly decreases with an increase in computing resources.
$\bullet$ Owing to the proposed gated self-attention module, HOGA is able to adaptively learn critical and high-order circuit structures, leading to $46.76\%$ error drop and $10.0\%$ accuracy improvement for QoR prediction and functional reasoning on new designs, respectively.
II Preliminary and Motivation
II-A Graph Neural Networks at Scale
GNNs have emerged as a promising technique that encodes circuit graphs into compact representations, which can then be utilized for addressing a wide spectrum of EDA problems. Specifically, given a circuit graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, GNNs update initial node features $\{x_{i}^{(0)}\;|\;i\in\mathcal{V}\}$ based on a node-wise aggregation scheme as follows:
$m_{i}^{(l)}=\text{AGGREGATE}^{(l)}(\{x_{j}^{(l-1)}\;|\;j\in\mathcal{N}(i)\})$ (1)
$x_{i}^{(l)}=\text{COMBINE}^{(l)}(m_{i}^{(l)},x_{i}^{(l-1)})$ (2)
where $x_{i}^{(l)}$ denotes the node feature vector at the $l$-th layer, $\mathcal{N}(i)$ represents the set of $1$-hop neighbors of node $i$, AGGREGATE can be any permutation-invariant operator such as MEAN or MAX pooling, and the goal of COMBINE function is to update node features based on the aggregated features from neighbors. After stacking $L$ GNN layers, the output features $\{x_{i}^{(L)}\;|\;i\in\mathcal{V}\}$ serve as the node representations that are used for downstream EDA applications.
However, producing a single node representation $x_{i}^{(L)}$ involves a recursive neighbor aggregation scheme, which can be straightforwardly derived by substituting Equation (1) into Equation (2). More concretely, let $d$ be the average node degree, there are $O(d^{L})$ nodes required to obtain $x_{i}^{(L)}$, which increases exponentially with the number of layers and is known as the “neighbor explosion” issue. Hence, GNNs generally suffer from poor scalability to large graphs, prohibiting their applications in large-scale circuit problems.
As a remedy, distributed training has been increasingly used for scaling GNNs to large graphs. However, handling the dependencies in a graph structure incurs significant communication overhead and work imbalance, hindering efficient distributed GNN training. To tackle this challenge, numerous efforts have been devoted to increase GNN training parallelism by adopting node-wise [9], layer-wise [4], and graph-wise sampling [20]. While these methods demonstrate promising results on common graph datasets such as social networks, we argue that they are unsuitable for circuit graphs, since the sampling algorithms may entirely break the design functionality and lead to poor accuracy. This is empirically confirmed in Section IV-C. In contrast, we introduce a novel hop-wise aggregation scheme, which independently learns each node representation based on its own hop-wise features. This renders our model embarrassingly parallel and greatly facilitates distributed training.
II-B Generalizable Graph Learning in EDA
Since realistic circuit designs may have different sizes or originate from distinct domains, generalization capability is crucial for deploying graph learning models in EDA applications. To this end, prior work has developed many customization techniques, particularly for GNNs in learning circuit representations. For instance, Ustun et al. proposed a customized GraphSAGE model by distinguishing predecessors and successors in a graph, which substantially outperforms the vanilla GraphSAGE on learning operation mapping patterns [15]. Later, Zhang et al. [21] and Guo et al. [8] introduced customized GNNs tailored for power inference and timing prediction tasks respectively, via sequentially updating node representations. More recently, Wu et al. presented a multi-task graph learning framework for functional reasoning. While these methods showcase promising results on the respective tasks, their GNN backbones are built upon the conventional message-passing paradigm, which cannot capture critical high-order structures formed by multiple nodes. It is worth noting that high-order structures are crucial for many circuit problems, such as functional reasoning. Specifically, identifying a half adder and a full adder within circuits requires to learn different high-order structures that correspond to their respective functional blocks [18]. Therefore, the aforementioned methods still cannot capture the intrinsic circuit information, leading to either restrictions to specific tasks or limited generalizability to complex circuit designs.
Although Wang et al. attempted to improve GNN generalizability by adopting the notion of graph contrastive learning [17], it involves nontrivial computational costs and thus worsens the GNN scalability. In this work, we aim to devise a graph learning model that is both scalable and generalizable for circuit designs, with the aid of a novel gated self-attention module applied on hop-wise features per node.
III The Proposed Approach
Figure 2: (a) An overview of the two major phases of HOGA; (b) Our introduced gated self-attention module for hop-wise feature aggregation.
Problem formulation. Given a circuit graph, we follow the common GNN settings that there are two inputs: graph adjacency matrix $A$ and node feature matrix $X$. Our goal is to build a model $\mathcal{M}$ for learning high-quality node representations $Y=\mathcal{M}(A,X)$, which can be utilized for a broad spectrum of circuit problems.
Figure 2(a) gives an overview of our proposed approach, HOGA, which consists of two major phases. During the first phase, HOGA computes hop-wise features by iteratively multiplying the adjacency matrix with the node feature matrix, as described in Section III-A. Note that this step can be finished in advance, allowing HOGA training with complexity independent of the graph structure. The second phase trains a novel gated self-attention module on hop-wise features to learn node representations. This module enables HOGA to adaptively learn critical and high-order features from different hops per node, as further illustrated in Section III-B. Consequently, HOGA is adaptive to different structures across various circuit designs while simultaneously benefiting from high parallelism for distributed training.
III-A Hop-Wise Feature Generation
Learning from neighborhood information is crucial for GNNs to capture critical graph structures. As depicted in Figure 1(b), traditional message-passing GNNs recursively aggregate features from neighbors, which leads to their poor scalability on large graphs. In contrast, we adopt a coarse-grained message-passing scheme based on hop-wise feature aggregation. To this ends, our first step is to generate hop-wise features.
Given a graph $\mathcal{G}$ with its adjacency matrix $A\in\mathbb{R}^{n\times n}$ and node feature matrix $X\in\mathbb{R}^{n\times d}$, where $n$ and $d$ represent the number of nodes and the feature dimension respectively, we first normalize the adjacency matrix: $\hat{A}=D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$, where $D$ is the node degree matrix. Next, we iteratively compute Equation (3) to generate hop-wise features for all nodes in parallel:
$X^{(k)}=\hat{A}X^{(k-1)},\;k=1,2,...,K$ (3)
where $X^{(0)}=X$, the $i$-th row of $X^{(k)}$ denotes the $k$-hop feature for node $i$, and $K$ is a hyperparameter to control the maximum number of hops for consideration.
After obtaining hop-wise features $X^{(0)},X^{(1)},...,X^{(K)}$, we stack them to construct a third-order tensor $\mathcal{X}\in\mathbb{R}^{n\times(K+1)\times d}$ such that:
$\mathcal{X}_{i}=[X_{i}^{(0)},X_{i}^{(1)},...,X_{i}^{(K)}]^{T},\;i=1,2,...,n$ (4)
Consequently, for each $i\in\{1,2,...,n\}$, $\mathcal{X}_{i}$ comprises up to $K$-hop features of node $i$, which are then independently used to learn the corresponding node representation $Y_{i}$. Hence, there are no dependencies between different nodes for learning their representations, making it easy to scale HOGA through distributed training. It is noteworthy that the time required for generating hop-wise features is generally negligible compared to the overall training time, as empirically confirmed in Section IV-B.
While there are a few prior arts (e.g. SIGN [7]) augmenting node features by adopting a similar approach to Equation (3), they simply train a multi-layer perceptron (MLP) model on augmented features. In contrast, we consider learning node representations through a hop-wise feature aggregation scheme, which is built upon a novel gated self-attention module introduced in the following section.
III-B Hop-Wise Aggregation with Gated Self-Attention
Since HOGA takes $\mathcal{X}_{i}\in\mathbb{R}^{(K+1)\times d}$ as input and produces a representation $Y_{i}\in\mathbb{R}^{d}$ for every node $i$ independently, we omit the node index $(i)$ in the ensuing discussion, by simply denoting $\mathcal{X}_{i}$ as $H$ and $Y_{i}$ as $y$ for clarity.
A straightforward way of producing $y$ is to accumulate hop-wise features in $H$, i.e., $y=\sum_{k=0}^{K}H_{k}$. However, this approach has two fundamental limitations: (1) it fails to capture high-order feature interactions among different hop neighbors, resulting in its incapability of learning high-order structures that are critical for many EDA problems (e.g., functional reasoning), as discussed in Section II-B. (2) it uniformly combines features from different hops and thus cannot identify and focus on important hop-wise features per node.
To address the aforementioned limitations, let us first consider a simple gated layer as follows:
$U=HW_{U},V=HW_{V}$ (5)
$\hat{H}=U\odot V$ (6)
where $W_{U},W_{V}\in\mathbb{R}^{d\times d}$ are trainable weight matrices, and $\odot$ denotes the element-wise product. By substituting Equation (5) into Equation (6), we can derive that $\hat{H}_{k}$ captures the second-order interaction $(H_{k}W_{U})\odot(H_{k}W_{V})$ for any hop $k\in\{0,1,...,K\}$. However, this also indicates that Equation (6) fails to capture interactions among different hop-wise features, i.e., $(H_{k}W_{U})\odot(H_{j}W_{V})$ with $k\neq j$. To tackle this limitation, we introduce a gated self-attention layer as follows:
$S=softmax(QK^{T})$ (7)
$\hat{H}=U\odot(SV)$ (8)
where $Q=HW_{Q},K=HW_{K}$, $W_{Q},W_{K}\in\mathbb{R}^{d\times d}$ are trainable weight matrices, and $S$ is the self-attention matrix widely used in Transformer-based models [16]. Through substituting Equation (5) into Equation (8), we can derive that $\hat{H}_{k}=(H_{k}W_{U})\odot(\sum_{j=0}^{K}S_{k,j}H_{j}W_{V})=\sum_{j=0}^{K}S_{k,j}\;(H_{k}W_{U})\odot(H_{j}W_{V})$, which captures second-order interactions on different hop-wise features. By stacking more layers, the output $\hat{H}$ naturally captures higher-order feature interactions from different hops, which correspond to higher-order structures on the input graph $\mathcal{G}$.
Figure 3: (a) An example of And-Inverter Graph, where each node represents a primary input/output (PI/PO) or an AND gate. The dashed arrows denote inverted edges; (b) An overview of QoR prediction. Our approach only replaces the GCN model in OpenABC-D with HOGA; (c) An overview of functional reasoning. Our approach solely substitutes HOGA for the GraphSAGE model in Gamora.
Implementation details. To improve the training stability of HOGA, we add LayerNorm and ReLU to Equation (8) in our implementation, resulting in the following:
$\hat{H}=\text{ReLU}(\text{LayerNorm}(U\odot(SV)))$ (9)
After obtaining $\hat{H}=[\hat{H}_{0},\hat{H}_{1},...,\hat{H}_{K}]^{T}\in\mathbb{R}^{(K+1)\times d}$, we exploit an attentive readout function as per Equation (10) to produce the final node representation $y$:
$\begin{split}c_{k}&=\frac{exp(\alpha^{T}(\hat{H}_{0}||\hat{H}_{k}))}{\sum_{j=1}^{K}exp(\alpha^{T}(\hat{H}_{0}||\hat{H}_{j}))}\\ y&=\hat{H}_{0}+\sum_{k=1}^{K}c_{k}\hat{H}_{k}\end{split}$ (10)
where $\alpha\in\mathbb{R}^{2d}$ is a trainable vector, ||||| | denotes the concatenation operator, and $c_{k}$ represents an attention score to measure the importance of the $k$-hop feature $\hat{H}_{k}$ to the final node representation $y$. In this way, HOGA can identify and adaptively aggregate critical features from different hops to produce high-quality node representations, which are then used for downstream tasks.
III-C Complexity analysis of HOGA
Given a graph with $n$ nodes and $m$ edges, suppose the feature dimension is $d\ll n$ and there are $K\ll n$ hops considered, the complexity of hop-wise feature generation is $O(Kmd)$. Besides, the gated self-attention module has the complexity of $O(nKd^{2}+nK^{2}d)$ for performing linear projections and computing the self-attention matrix $S$ in Equation (7). Thus, the total complexity of HOGA is $O(Kmd+nKd^{2}+nK^{2}d)$, which is linear with respect to the number of nodes/edges.
IV Experiment
Table 1: Statistics of OpenABC-D benchmark — The upper $20$ designs are used for training and the lower $9$ designs for evaluation.
IP Design Nodes Edges Category
spi $4219$ $8676$ Communication
i2c $1169$ $2466$ Communication
ss_pcm $462$ $896$ Communication
usb_phy $487$ $1064$ Communication
sasc $613$ $1351$ Communication
wb_dma $4587$ $9876$ Communication
simple_spi $930$ $1992$ Communication
pci $19547$ $42251$ Communication
dynamic_node $18094$ $38763$ Control
ac97_ctrl $11464$ $25065$ Control
mem_ctrl $16307$ $37146$ Control
des3_area $4971$ $10006$ Crypto
aes $28925$ $58379$ Crypto
sha256 $15816$ $32674$ Crypto
fir $4558$ $9467$ DSP
iir $6978$ $14397$ DSP
idft $241552$ $520523$ DSP
dft $245046$ $527509$ DSP
tv80 $11328$ $23017$ Processor
fpu $29623$ $59655$ Processor
wb_conmax $47840$ $97755$ Communication
ethernet $67164$ $144750$ Communication
bp_be $82514$ $173441$ Control
vga_lcd $105334$ $227731$ Control
aes_xcrypt $45840$ $93485$ Crypto
aes_secworks $40778$ $84160$ Crypto
jpeg $114771$ $234331$ DSP
tiny_rocket $52315$ $108811$ Processor
picosoc $82945$ $176687$ Processor
IV-A Experimental Setup
Table 2: Comparison of HOGA and GCN for QoR prediction — We choose mean absolute percentage error (MAPE) as the evaluation metric. Lower MAPE score means better QoR prediction. HOGA-$2$ and HOGA-$5$ indicate $K=2$ and $K=5$ in HOGA, respectively.
wb_conmax ethernet bp_be vga_lcd aes_xcrypt aes_secworks jpeg tiny_rocket picosoc Average Training Time GCN $24.66\%$ $19.20\%$ $39.53\%$ $52.35\%$ $21.41\%$ $23.93\%$ $22.11\%$ $19.89\%$ $10.51\%$ $26.0\%$ $11.9$ hours (1.0×1.0$\times$1.0 ×) HOGA-$2$ $7.33\%$ $6.51\%$ $17.81\%$ $9.31\%$ $\mathbf{5.79\%}$ $10.17\%$ $\mathbf{3.87\%}$ $6.70\%$ $2.91\%$ $7.8\%$ $\mathbf{3.8}$ hours (3.1$\downarrow$ 3.1$\times$ 3.1 ×) HOGA-$5$ $\mathbf{4.53\%}$ $\mathbf{4.21\%}$ $\mathbf{4.76\%}$ $\mathbf{5.59\%}$ $6.57\%$ $\mathbf{8.42\%}$ $5.65\%$ $\mathbf{3.88\%}$ $\mathbf{1.90\%}$ $\mathbf{5.0\%}$ $11.2$ hours (1.1$\downarrow$ 1.1$\times$ 1.1 ×)
Figure 4: GCN and HOGA-$5$ predictions vs. ground truth — The ground truth value is preprocessed in the OpenABC-D benchmark.
We evaluate HOGA on two representative EDA tasks: (1) QoR prediction. We focus on the OpenABC-D benchmark for predicting the optimized number of gates in And-Inverter-Graphs (AIGs) after logic synthesis optimization through ABC [3]. Notably, OpenABC-D consists of $870,000$ AIGs that are generated by running various synthesis recipes on $29$ open-source hardware IPs shown in Table 1, which contain different functionality designs from MIT LL labs CEP [5], OpenCores [1], and IWLS [2]. To demonstrate the generalizability of our approach, we train HOGA on the upper $20$ designs in Table 1 and evaluate it on the rest of the designs, which is a challenging graph-level regression task; (2) Functional reasoning. We follow Gamora [18] to identify adder blocks by predicting the sum and carry-out nodes in AIGs of carry-save array (CSA) and Booth multipliers, which is formulated as a node-level classification task. As is known that technology mapping can largely increase the reasoning complexity on Boolean networks [12, 19], we consider the most challenging scenario in Gamora, where AIGs are generated by ABC with complex ASAP 7nm technology mapping, and we only train HOGA on an $8$-bit multiplier design and perform inference on multipliers with bitwidth up to $768$.
Note that the authors of OpenABC-D and Gamora have proposed their own GNN-based methods for the aforementioned tasks respectively, which serve as our baselines. To ensure a fair comparison, we only replace their GNN models with HOGA and keep other components the same in both baseline methods, as depicted in Figure 3. In regard to hyperparameter settings of HOGA, we adopt the Adam optimizer with a learning rate of $0.0001$, a hidden dimension of $256$, and fix the number of gated self-attention layer to $1$. Besides, we set the number of hops $K$ as $5$ for experiments on OpenABC-D and $8$ on Gamora, which captures the information from the same number of hops as the baseline GNNs in both tasks. We leverage DistributedDataParallel in the PyTorch framework for distributed training on HOGA. All experiments are conducted on a Linux machine with an Intel Xeon Gold $5218$ CPU, which features 8 cores running at $2.30$ GHz, and is equipped with 4 RTX A6000600060006000 GPUs, each having 48 GB of memory.
IV-B Evaluation on QoR Prediction
Figure 5: Multi-GPU training time of HOGA on OpenABC-D.
Figure 6: Functional reasoning on AIG-based CSA and Booth multipliers with $7$nm technology mapping.
Our baseline is a $5$-layer GCN, as previously used for the OpenABC-D benchmark [6].
Figure 7: Visualization on hop-wise attention scores per node in the AIG-based $768$-bit Booth Multiplier — Rows and columns in each heatmap denote nodes and their hop-wise neighbors, respectively.
Besides, we choose mean absolute percentage error (MAPE) as the evaluation metric, which is defined as $\text{MAPE}=\frac{1}{g}\sum_{i=1}^{g}|\frac{y_{i}-\hat{y}_{i}}{y_{i}}|\times 100\%$, where $y_{i}$ and $\hat{y}_{i}$ denote the ground truth and model prediction on the $i$-th sample (graph), respectively. Table 2 indicates that both HOGA models significantly outperform GCN across all test designs. In particular, HOGA-$5$ improves the estimation error over GCN on vga_lcd by a margin of $46.76\%$. The superiority of HOGA is further demonstrated in Figure 4, where the QoR predictions by HOGA-$5$ are highly correlated with the ground truth. In contrast, the GCN model fails to accurately predict the actual QoR values on those unseen designs. Moreover, by comparing HOGA-$5$ and HOGA-$2$, we can see the trade-off between the accuracy and training time of HOGA. Compared to the GCN baseline, HOGA-$2$ not only improves the average estimation error by $18.2\%$, but also achieves a notable 3.1×3.1$\times$3.1 × training speedup on a single GPU.
It is noteworthy that HOGA can be easily accelerated through distributed training, owing to its high parallelism for learning node representations. As shown in Figure 5, the training time of HOGA almost linearly decreases when the number of GPU devices is increased. As a result, it takes only $3.5$ ($1.1$) hours to train HOGA-$5$ (HOGA-$2$) on the OpenABC-D benchmark. Notably, the first phase of HOGA for generating hop-wise features takes $13$ minutes, which is negligible compared to the overall training time. We believe that HOGA can be further accelerated if more computing resources are available, rendering it applicable to industrial-scale circuit applications.
IV-C Evaluation on Functional Reasoning
To further demonstrate the generalization capability, we evaluate HOGA on the functional reasoning task [18] by identifying root nodes of XOR and majority (MAJ) operations in AIG-based multipliers, which correspond to sum and carry-out of adder blocks, respectively. More concretely, there are $4$ node categories for classification: (1) MAJ nodes; (2) XOR nodes; (3) nodes shared by MAJ and XOR; (4) other plain nodes such as PI/PO. Apart from choosing the GraphSAGE model used in Gamora as our baseline, we also consider GraphSAINT [20], a popular sampling-based GNN model, and SIGN [7] that adopts an MLP model on hop-wise features. Note that all baselines as well as HOGA are trained on an AIG-based $8$-bit multiplier and evaluated on multipliers with bitwidths ranging from $64$ to $768$.
As shown in Figure 6, all baselines perform poorly on CSA multipliers, which indicates their limited generalizability to complex designs. In addition, GraphSAINT performs even worse than the vanilla GraphSAGE model on Booth multipliers, indicating sampling-based GNNs are not suitable for circuit graphs, as analyzed in Section II-A. While SIGN achieves reasonable accuracy on Booth multipliers, it still largely lags behind HOGA on CSA multipliers. We attribute it to our proposed gated self-attention module, which adaptively learns high-order features from different hops per node. This is further confirmed in Section IV-D. As a result, HOGA largely surpasses baseline models on both CSA and Booth multiplier circuits, with a $10.0\%$ accuracy improvement on the $768$-bit CSA multiplier. More importantly, Figure 6 illustrates that the accuracy of HOGA exhibits a rising trend as the bitwidth of test multipliers increases, which is crucial for accurate reasoning on large-scale Boolean networks.
IV-D Visualization on HOGA Attention Scores
For the sake of clear visualization, we randomly sample $100$ nodes per classification category from the AIG of $768$-bit Booth multiplier, based on which four heatmaps are drawn in Figure 7. Notably, the rows and columns in each heatmap correspond to nodes and their hop-wise neighbors, respectively. Each element in a heatmap row represents the hop-wise attention score $c_{k}$ defined in Equation (10), which is used to adaptively combine $\hat{H}_{k}$ for producing the final node representations. Figure 7 clearly shows that HOGA is able to identify critical features from different hop $k$ for learning the underlying Boolean functions. Notably, since we use a single gated self-attention layer, each $\hat{H}_{k}$ captures second-order interactions among hop-wise features, corresponding to second-order graph structures. As a consequence, HOGA skips odd-hop neighbor features and primarily focuses on $\hat{H}_{k}$ with $k\in\{2,4,6\}$ for learning MAJ/XOR functions, as indicated by the attention scores shown in the first three heatmaps. As for the last heatmap, the attention scores of HOGA are relatively random since those nodes are PI/PO or plain AND gates, whose Boolean functions are not meaningful for this task.
V Conclusion
This work introduces HOGA, a hop-wise attention approach for learning circuit representations in a scalable and generalizable way. HOGA first computes hop-wise features, which are then independently fed into a gated self-attention module for capturing high-order interactions among different hop neighbors. Experimental results showcase that HOGA not only comfortably scales to large-scale circuits via distributed training, but also outperforms conventional GNNs for generalizing to unseen and complex circuit designs.
Acknowledgments
This work is supported in part by a Qualcomm Innovation Fellowship, NSF Awards #2118709 and #2212371, and ACE, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.